Find the value of x. Leave your answer in simplest radical form.
Piture is a right trianggle and base is an x which is the bottom and then on the left side of the right triangle is 8 and the right side of the right trangle is 14.

Answer by stanbon(75887)

Find the value of x. Leave your answer in simplest radical form.
Piture is a right trianggle and base is an x which is the bottom and then on the left side of the right triangle is 8 and the right side of the right trangle is 14.
-----------------------
IF x is the hypotenuse and 8 and 14 are the lengths of the legs,
x^2 = 4^2 + 14^2
x^2 = 212
x^2 = 4*53
x = 2*sqrt(53)
===============
Cheers,
Stan H.